# VoiceCloak: A Multi-Dimensional Defense Framework against Unauthorized Diffusion-based Voice Cloning

### v4

@@ 1 Introduction
Diffusion Models (DMs) (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020; Rombach et al. 2022) have recently emerged as powerful generative tools, achieving unprecedented success within realistic voice cloning (VC). Their iterative denoising process enables generating speech with remarkable naturalness, detail, and fidelity to human voice (Popov et al. 2022; Kong et al. 2020; Jeong et al. 2021; Shen et al. 2024). However, the open-source availability and ease of use of these models intensify concerns about potential misuse. Attackers can synthesize highly realistic voice replicas from short public audio clips, as depicted in Figure 1(a), enabling sophisticated fraud and circumvention of voiceprint authentication.
@@
To counter such unauthorized use, two main defense paradigms are introduced, including forgery detection and proactive disruption. Reactive detection methods (Jung et al. 2022; Zhou and Lim 2021; Wu et al. 2024) identify forgeries after they are crafted, often too late to prevent harm. This highlights the need for proactive defenses that disrupt the synthesis process itself. Prior proactive work (Huang et al. 2021; Yu, Zhai, and Zhang 2023; Chen et al. 2024; Li et al. 2023) has focused on adding imperceptible adversarial perturbations to reference audio by compromising the functionality of either the voice decoder or the speaker identity encoder.
However, existing defenses designed for prior architectures are largely ineffective against Diffusion Models (DMs). This incompatibility arises from two fundamental challenges: gradient vanishing and dynamic conditioning (Figure 1). Specifically, (1) the single forward pass gradient computation relied upon by many defenses become unreliable or vanish within the multi-step denoising process of DMs and the corresponding deep computational graph, rendering such single-pass gradient information ineffective for disrupting the full generation trajectory (Kang, Song, and Li 2023). (2) Strategies targeting specific subnetworks (e.g. speaker or content encoders) fail because DMs often employ dynamic conditioning mechanisms, which means no single modules solely responsible for condition processing. Consequently, methods targeting individual components struggle to cause global disruption. These fundamental incompatibilities underscore the need for novel strategies tailored to this generative paradigm.
@@ 2.1 Audio Diffusion Models
Diffusion Models (Ruan et al. 2023; Kawar et al. 2023; Ruiz et al. 2023) a dominant force in generative modeling, demonstrating extraordinary performance across multi-modal tasks, significantly advancing audio synthesis tasks like text-to-speech (Popov et al. 2021; Jeong et al. 2021) and unconditional audio generation (Kong et al. 2020; Liu et al. 2023). Particularly within voice cloning (VC), diffusion-based methods(Popov et al. 2022; Choi, Lee, and Lee 2024), mostly leveraging score-based formulations via stochastic differential equations (SDEs) (Song et al. 2020), now yield outputs with remarkable naturalness and speaker fidelity. While impressive, this state-of-the-art performance significantly heightens concerns regarding potential misuse, directly motivating proactive defense strategies such as the framework proposed herein.
@@ 2.2 Proactive Defense via Adversarial Examples
Beyond passive DeepFake detection, proactive defenses aim to preemptively disrupt malicious syntheses, by introducing adversarial perturbations to the original audio. Early work (Huang et al. 2021) demonstrated the feasibility of this approach but struggled to balance effectiveness with imperceptibility. Subsequent research focused on improving this trade-off. Strategies included using psychoacoustic models to enhance imperceptibility (Li et al. 2023), incorporating human-in-the-loop refinement for better balance (Yu, Zhai, and Zhang 2023), , and employing GAN-based generators to improve efficiency (Dong et al. 2024).
@@ 3.1 Score-based Diffusion
Score-based generative models define a continuous-time diffusion process using stochastic differential equations (SDEs) (Song et al. 2020). The forward process gradually perturbs clean data $x_{0}\sim p_{0}(x)$ into noise via the SDE:
@@
where $f(\mathbf{x},t)$ and $g(t)$ are the drift and diffusion coefficients, and $\mathbf{w}$ is a Wiener process. The corresponding reverse-time SDE that transforms $x_{T}$ back into $p_{0}(x)$ can be expressed as:
@@ 3.2 Adversarial Vulnerability Analysis
Additionally, we target the U-Net’s internal feature representations as a vulnerability for degrading perceptual quality. Prior attribute editing work (Ronneberger, Fischer, and Brox 2015; Oh, Lee, and Lee 2024; Tumanyan et al. 2023) confirms that these hierarchical features within the U-Net are controllable and also encode crucial semantic and acoustic details that govern the coherence and naturalness of the synthesized speech. Therefore, adversarially corrupting these representations directly impair the model’s ability to synthesize natural-sounding speech, offering a complementary strategy for fidelity degradation. This analysis reveals critical vulnerabilities in DMs, providing targeted avenues for adversarial intervention.
@@ 4.2 Adversarial Identity Obfuscation
Psychoacoustic studies (Kreiman and Sidtis 2011; Lavner, Gath, and Rosenhouse 2000) indicate that significant perceptual differences in speaker identity often exist between genders, linked to specific acoustic cues like F0 and formant structures . Leveraging this, we assume that guiding the adversarial embedding towards the opposite gender one will likely create the strongest perceptual contrast, enhancing identity obfuscation.
@@
$\mathcal{L}_{ID}=-Sim(\mathcal{R}_{adv},\mathcal{R}_{ref}+\underbrace{Sim(\mathcal{R}_{adv},\mathcal{C}_{opp})}_{\text{Gender}},$ (7)
@@
Commonly in diffusion-based VC, conditional information from $x_{ref}$ is integrated via Linear-attention layers (Katharopoulos et al. 2020). Within the U-Net, latent code representing the content $x_{src}^{t}$ are linearly projected to $Q$ matrix, whereas condition latents $x_{ref}^{t}$ are projected to $K$ and $V$. Linear-attention first computes a context matrix by aggregating values ($V$) weighted by keys ($K$):
$\mathcal{H}_{ctx}(x_{ref})=\text{Softmax}(\phi^{(l,t)}(x_{ref}^{t})W_{K}^{l})(\phi^{(l,t)}(x_{ref}^{t})W_{V}^{l})^{T},$
where $\mathcal{H}_{ctx}$ represents the context hidden state, $\phi^{(l,t)}(\cdot)$ is the deep features of the $l^{th}$ block in U-Net at timestep $t$, and $W_{K}^{l},W_{V}^{l}$ are the projection matrices. Then, the context interacts with queries ($Q$) to obtain the final attention output: $\mathcal{A}^{l}=(\mathcal{H}_{ctx})^{T}(W_{Q}^{l}\phi^{(l,t)}(x_{src}^{t}))$. This context representation provides a dynamic summary of the reference speaker’s stylistic features, weighted by their relevance to the current content queries during synthesis. By applying a softmax function, we obtain an explicit probability distribution over sequence positions which signifies the model’s information focus. Our strategy, therefore, is to maximize the Kullback-Leibler (KL) divergence (Kullback and Leibler 1951) between the context distribution derived from the original reference and the adversarial audio:
@@ 5.1 Experimental Setup
Experiments are conducted on the LibriTTS (Zen et al. 2019) and VCTK (Yamagishi, Veaux, and MacDonald 2019) datasets. We selected a gender-balanced audio subsets (479 utterances from LibriTTS, 500 from VCTK) to generate adversarial reference speech set $\mathcal{D}_{x_{adv}}$.
@@ 5.4 Transferability
To demonstrate applicability, we extended our experiments to include two additional open-source diffusion-based VC models: DuTa-VC (Wang et al. 2023) and DDDM-VC (Choi, Lee, and Lee 2024). As shown in Table 4, VoiceCloak demonstrate favorable transferability to different models, achieving an average DSR of 66.7%. We attribute this transferability to: Targeting Common Vulnerabilities, as our method exploits fundamental mechanisms (e.g. attention, score prediction) that are shared across diffusion VCs.

### v5

@@ 1 Introduction
Diffusion Models (DMs) (Ho et al. 2020; Song et al. 2020a; Rombach et al. 2022) have recently emerged as powerful generative tools, achieving unprecedented success within realistic voice cloning (VC). Their iterative denoising process enables generating speech with remarkable naturalness, detail, and fidelity to human voice (Popov et al. 2022; Kong et al. 2020; Jeong et al. 2021; Shen et al. 2024). However, the open-source availability and ease of use of these models intensify concerns about potential misuse. Attackers can synthesize highly realistic voice replicas from short public audio clips, as depicted in Figure 1(a), enabling sophisticated fraud and circumvention of voiceprint authentication.
@@
To counter such unauthorized use, two main defense paradigms are introduced, including forgery detection and proactive disruption. Reactive detection methods (Jung et al. 2022; Zhou and Lim 2021; Wu et al. 2024) identify forgeries after they are crafted, often too late to prevent harm. This highlights the need for proactive defenses that disrupt the synthesis process itself. Prior proactive work (Huang et al. 2021; Yu et al. 2023; Chen et al. 2024; Li et al. 2023) has focused on adding imperceptible adversarial perturbations to reference audio by compromising the functionality of either the voice decoder or the speaker identity encoder.
However, existing defenses designed for prior architectures are largely ineffective against Diffusion Models (DMs). This incompatibility arises from two fundamental challenges: gradient vanishing and dynamic conditioning (Figure 1). Specifically, (1) the single forward pass gradient computation relied upon by many defenses become unreliable or vanish within the multi-step denoising process of DMs and the corresponding deep computational graph, rendering such single-pass gradient information ineffective for disrupting the full generation trajectory (Kang et al. 2023). (2) Strategies targeting specific subnetworks (e.g. speaker or content encoders) fail because DMs often employ dynamic conditioning mechanisms, which means no single modules solely responsible for condition processing. Consequently, methods targeting individual components struggle to cause global disruption. These fundamental incompatibilities underscore the need for novel strategies tailored to this generative paradigm.
@@ 2.1 Audio Diffusion Models
Diffusion Models (Ruan et al. 2023; Kawar et al. 2023; Ruiz et al. 2023) a dominant force in generative modeling, demonstrating extraordinary performance across multi-modal tasks, significantly advancing audio synthesis tasks like text-to-speech (Popov et al. 2021; Jeong et al. 2021) and unconditional audio generation (Kong et al. 2020; Liu et al. 2023). Particularly within voice cloning (VC), diffusion-based methods(Popov et al. 2022; Choi et al. 2024), mostly leveraging score-based formulations via stochastic differential equations (SDEs) (Song et al. 2020b), now yield outputs with remarkable naturalness and speaker fidelity. While impressive, this state-of-the-art performance significantly heightens concerns regarding potential misuse, directly motivating proactive defense strategies such as the framework proposed herein.
@@ 2.2 Proactive Defense via Adversarial Examples
Beyond passive DeepFake detection, proactive defenses aim to preemptively disrupt malicious syntheses, by introducing adversarial perturbations to the original audio. Early work (Huang et al. 2021) demonstrated the feasibility of this approach but struggled to balance effectiveness with imperceptibility. Subsequent research focused on improving this trade-off. Strategies included using psychoacoustic models to enhance imperceptibility (Li et al. 2023), incorporating human-in-the-loop refinement for better balance (Yu et al. 2023), , and employing GAN-based generators to improve efficiency (Dong et al. 2024).
@@ 3.1 Score-based Diffusion
Score-based generative models define a continuous-time diffusion process using stochastic differential equations (SDEs) (Song et al. 2020b). The forward process gradually perturbs clean data $\mathbf{x}_{0}\sim p_{0}(\mathbf{x})$ into noise via the SDE:
@@
where $f(\mathbf{x},t)$ and $g(t)$ are the drift and diffusion coefficients, and $\mathbf{w}$ is a Wiener process. The corresponding reverse-time SDE that transforms $x_{T}$ back into $p_{0}(\mathbf{x})$ can be expressed as:
@@ 3.2 Adversarial Vulnerability Analysis
Additionally, we target the U-Net’s internal feature representations as a vulnerability for degrading perceptual quality. Prior attribute editing work (Ronneberger et al. 2015; Oh et al. 2024; Tumanyan et al. 2023) confirms that these hierarchical features within the U-Net are controllable and also encode crucial semantic and acoustic details that govern the coherence and naturalness of the synthesized speech. Therefore, adversarially corrupting these representations directly impair the model’s ability to synthesize natural-sounding speech, offering a complementary strategy for fidelity degradation. This analysis reveals critical vulnerabilities in DMs, providing targeted avenues for adversarial intervention.
@@ 4.2 Adversarial Identity Obfuscation
Psychoacoustic studies (Kreiman and Sidtis 2011; Lavner et al. 2000) indicate that significant perceptual differences in speaker identity often exist between genders, linked to specific acoustic cues like F0 and formant structures . Leveraging this, we assume that guiding the adversarial embedding towards the opposite gender one will likely create the strongest perceptual contrast, enhancing identity obfuscation.
@@
$\mathcal{L}_{ID}=-Sim(\mathcal{R}_{adv},\mathcal{R}_{ref})+\underbrace{Sim(\mathcal{R}_{adv},\mathcal{C}_{opp})}_{\text{Gender}},$ (7)
@@
Commonly in diffusion-based VC, conditional information from $x_{ref}$ is integrated via Linear-attention layers (Katharopoulos et al. 2020). Within the U-Net, latent code representing the content $x_{src}^{t}$ are linearly projected to $Q$ matrix, whereas condition latents $x_{ref}^{t}$ are projected to $K$ and $V$. Linear-attention first computes a context matrix by aggregating values ($V$) weighted by keys ($K$): H_ctx(x_ref)=Softmax(ϕ^(l,t)(x_ref^t) W_K^l)(ϕ^(l,t)(x_ref^t) W_V^l)^T, where $\mathcal{H}_{ctx}$ represents the context hidden state, $\phi^{(l,t)}(\cdot)$ is the deep features of the $l^{th}$ block in U-Net at timestep $t$, and $W_{K}^{l},W_{V}^{l}$ are the projection matrices. Then, the context interacts with queries ($Q$) to obtain the final attention output: $\mathcal{A}^{l}=(\mathcal{H}_{ctx})^{T}(W_{Q}^{l}\phi^{(l,t)}(x_{src}^{t}))$. This context representation provides a dynamic summary of the reference speaker’s stylistic features, weighted by their relevance to the current content queries during synthesis. By applying a softmax function, we obtain an explicit probability distribution over sequence positions which signifies the model’s information focus. Our strategy, therefore, is to maximize the Kullback-Leibler (KL) divergence (Kullback and Leibler 1951) between the context distribution derived from the original reference and the adversarial audio:
@@ 5.1 Experimental Setup
Experiments are conducted on the LibriTTS (Zen et al. 2019) and VCTK (Yamagishi et al. 2019) datasets. We selected a gender-balanced audio subsets (479 utterances from LibriTTS, 500 from VCTK) to generate adversarial reference speech set $\mathcal{D}_{x_{adv}}$.
@@ 5.4 Transferability
To demonstrate applicability, we extended our experiments to include two additional open-source diffusion-based VC models: DuTa-VC (Wang et al. 2023) and DDDM-VC (Choi et al. 2024). As shown in Table 4, VoiceCloak demonstrate favorable transferability to different models, achieving an average DSR of 66.7%. We attribute this transferability to: Targeting Common Vulnerabilities, as our method exploits fundamental mechanisms (e.g. attention, score prediction) that are shared across diffusion VCs.